OpenStudy (anonymous):

x+y=4 solving by substition x-y=8.8

OpenStudy (anonymous):

Add the two equations together: 2x=12.8 => x=6.4 6.4+y=4=>y=-2.4 6.4-(-2.4)=>8.8 as required

OpenStudy (anonymous):

x + y = 4 x= 4- y........(i) x - y = 8.8.....(ii) Substituting i in ii 4 - y - y = 8.8 4 - 2y = 8.8.......... -2y = 8.8 - 4 -2y = 4.8 devide both sides by -2 y= -2.4 then substitute y in i......... x = 4 - -2.4 x= 6.4 can confirm this with equation i 6.4 + (- 2.4) 6.4 - 2.4 = 4